# Perfect Superconducting Diode and Supercurrent Range Controller

### v1

@@ Abstract
Diodes have a nonreciprocal voltage versus current relationship, produced by breaking the space and time reversal symmetry. However, developing high-end superconducting computers requires a superconducting analogue of the traditional semiconductor diode. Such a superconducting diode exhibits non-reciprocity, or a high asymmetry in its critical currents. We present a model of a perfect superconducting diode based on a superconducting quantum interference device made with multiple superconducting nanowires (MW-SQUID). The superconducting condensate in the nanowires is assumed to obey a linear current-phase relationship (CPR). The diode predicted by our model has a large positive critical current, while the negative critical current can be made exactly zero. Thus achieved 100% diode efficiency ($\eta=1$) remains stable against small changes of the magnetic field. Another important result is that under certain and quite broad conditions such devices can act as supercurrent range controllers. In such device a supercurrent can flow with zero voltage applied, but only if the supercurrent is contained in some narrow, adjustable range, which excludes zero current.
@@
Superconducting electronics offer a significant energy advantage over that of their traditional semiconductor counterpart due the absence of resistivity Tinkham (2004). This drastically improves energy scalability of superconducting computing resources.
@@
Superconducting diodes are useful for quantum and cryogenic electronics with their ability to allow supercurrent to flow preferentially in one direction Souto, Leijnse, and Schrade (2022); Chen et al. (2018); Hoshino et al. (2018); Baumgartner et al. (2021). In quantum computing, superconducting diodes are crucial for efficient power delivery in quantum processors, ensuring low-noise operation at ultra-cold temperatures Ingla-Aynés et al. (2025); Ideue and Iwasa (2020). In cryogenic electronics, superconducting diodes enable rectification circuits that convert AC to DC in cryogenic temperatures which improves energy efficiency Tuhin (2025). Unique properties of superconducting diodes make them ideal for ultra-sensitive detection systems and sensors used in medical imaging Ando et al. (2020); Hou et al. (2023). Superconducting diodes might play a role in advanced superconducting circuits and enable the development of the next-generation superconducting logic devices and memory elements Ma, Zhan, and Lin (2025); Alam et al. (2021); Ilin et al. (2021); Song et al. (2023). Note that an intrinsic perfect superconducting diode (PSD) has been discussed theoretically Chakraborty and Black-Schaffer (2024). Additionally, Josephson Junctions have been used, but are limited due to their large size Holmes (2021); Herr, Josephsen, and Herr (2023). Therefore, new approaches are necessaryGolod, Morlet-Decarnin, and Krasnov (2023). We propose an alternative approach to construct PSD using superconducting nanowires which doesn’t require any exotic superconductor and can be made out of ordinary superconducting metals such as niobium or aluminum.
@@
We present a simple yet efficient model of a superconducting quantum interference device (SQUID) made with a few nanowires acting as parallel weak links. One key difference with respect to the ordinary SQUID is that our nanowires are assumed to have a linear current-phase relationship, as established previouslyMurphy and Bezryadin (2017); Murphy, Averin, and Bezryadin (2017); Ilin et al. (2021). The phase difference $\phi$ is limited by the critical phase $\phi_{c}$, proportional to the wire length. The linear CPR means that the supercurrent, $I$, is proportional to the phase gradient of the condensate wave function (similar to free quantum particle). The phase gradient, on the other hand, is proportional to the phase difference $\phi$ between the ends of the nanowire. The linear CPR enables the manifestation of multiple highly stable vorticity states. Our modelSun and Bezryadin (2024) allows us to calculate vorticity stability regions (VSR), which are the regions in the current-magnetic (I-B) field plane within which the total supercurrent is within the negative and positive critical current. We discover that some of the vorticity states generate VSRs that act as an supercurrent range controller (SRC) or as a perfect superconducting diode (PSD). In an SRC, the supercurrent ($I$) is confined between two positive values, i.e. $0<I_{c,-}<I<I_{c,+}$. In a PSD, the supercurrent is confined as $0=I_{c,-}<I<I_{c,+}$, which corresponds to the perfect efficiency $\eta=(I_{c,+}+I_{c,-})/(I_{c,+}-I_{c,-})=1$. Note that since in the SRC $I_{c,-}>0$, therefore formally speaking $\eta>1$.
@@
In this section, we discuss our computational model. We consider a parallel array of superconducting nanowires connected to two macroscopic thin-film superconducting electrodes. To model this device, we assume that all the superconducting nanowires obey a linearTinkham (2004) current-phase-relationship (CPR) as given by:
@@
where $i$ is the wire number, $\phi_{i}$ is the phase difference across the wire $i$, and $\phi_{c,i}$ and and $I_{c,i}$ represent the critical phase and the critical current of the wire $i$. We assume that the entire device switches to the normal (non-superconducting) state as soon as any single wire switches, which happens if $\phi_{i}=\phi_{c,i}$ for any $i$. Thermal and quantum phase fluctuations are assumed to be negligible.
@@
Using Eqs.(3),(4) we generate vorticity stability regions (VSR), each defined as region on the $I-B$ plane where for a fixed vorticity, the current in each wire is lower than its corresponding critical current. The boundary of the VSR is the critical current of the device as a function of the magnetic field. If either the total current, $I$, or the magnetic field, $B$ is changed beyond the VSR boundaries, then the device switches to the normal (i.e., resistive) state.
@@
Figure 2: A SRC is demonstrated for VSRs generated by (a) a 2-SQUID with critical phases set to $[\pi,2\pi]$ and critical currents set to $[2/3,4/3]$. The plotted vortices are $-1$ (red), 00 (blue), and $1$ (black). And (b) a disordered 3-SQUID with wires at $[0,0.45,1]$, critical phases set to $[6,7,5]$, and critical currents set to $[1.01,1.12,0.87]$. The vortices from left to right are $[2,-1]$ (red), $[3,0]$ (black), $[4,1]$ (blue). The grey lines indicate any other VSRs corresponding to different vorticity states.
Let us first construct an example of a point-wise PSD. Suppose the SQUID has only two wires, $n=2$, with $v_{1,2}$ vortices trapped between them (see Fig.1). Then the MPC equation is $\phi_{2}=\phi_{1}+2\pi b-2\pi v_{1,2}$. Consider the case $v_{1,2}=0$, so $\phi_{2}=\phi_{1}+2\pi b$. Zero critical current means that even if the bias current is zero the magnitude of the current in one of the wire is equal to its critical current. Therefore, either $\phi_{1}=\phi_{c,1}$ or $\phi_{1}=-\phi_{c,1}$. Suppose $\phi_{1}=-\phi_{c,1}$. Then $\phi_{2}=-\phi_{c,1}+2\pi b$. Under such condition one can increase the bias current and the device will remain superconducting. But a decrease of the bias current is not possible since the first wire would be pushed into the normal state. To construct a PSD, require that the total current is zero, $I=0$, and that the first wire is in the critical state. If $I=0$ then the current in the first wire is equal in magnitude and opposite in sign to the current in the second wire, i.e. $\phi_{1}I_{c,1}/\phi_{c,1}=-\phi_{2}I_{c,2}/\phi_{c,2}$. If $\phi_{1}=-\phi_{c,1}$ then $-\phi_{c,1}I_{c,1}/\phi_{c,1}=-(-\phi_{c,1}+2\pi b)I_{c,2}/\phi_{c,2}$. So $I_{c,1}=(-\phi_{c,1}+2\pi b)I_{c,2}/\phi_{c,2}$. For this to be true the right hand side needs to be positive, because $I_{c,1}$ represents the magnitude of the critical current of the first wire. As an example, choose $b=1$ and $\phi_{c,1}=\pi$. Then $I_{c,1}=\pi I_{c,2}/\phi_{c,2}$. Suppose $\phi_{c,2}=2\pi$. Then $I_{c,1}=I_{c,2}/2$. This argument indicates that a 2-SQUID with critical phases $[\pi,2\pi]$ and critical currents, for example, $[2/3,4/3]$ should have a zero critical current at the point $b=1$, i.e., it should act as a perfect superconducting diode. This device will be called point-wise PSD since $I_{c,-}(b)=0$ is only true if $b=1$, or other integers if $v_{1,2}\neq 0$ (Fig.2a).
If $v_{1,2}=0$ and if $b>1$, then the allowed range of the supercurrent becomes strictly above zero (Fig.2a). For example, if $b=1.25$ then the allowed range of the supercurrent is $10.3<I<$ 1. Thus at non-zero magnetic fields our PSD becomes a supercurrent range controller (SRC). The VSR repeats with the period $\Delta b=1$, due to the Little-Parks effect in this 2-SQUID.
@@
For the critical current of one polarity to be independent of the magnetic field, it must be the case that the $b$ term drops out. In other words, $02\pi b\sum_{i=1}^{n}(x_{i}-x_{m})=$ 0 or:
@@
Figure 3: A b-invariant PSD can be constructed by (a) a symmetrical 3-SQUID with $\phi_{c}=2\pi$ for all wires, setting vorticity states [1,-2] (red) and [2, -1] (blue) and by (b) a non-symmetrical 5-SQUID with wires at $[0,5/8,2/3,5/6,1]$ and critical phases set to $512\pi/$ 5. The vorticity state is $[3,-1,0,0]$. In both figures, the underlying gray lines indicate other VSRs corresponding to different vorticity states.
@@
Next, we construct an example of a b-invariant PSD in a symmetrical 3-SQUID. Consider a 3-SQUID device with all critical phases set to $2\pi$. According to Eq.(7), to produce a VSR with the flat and horizontal bottom one needs $\sum_{i=1}^{n}v_{i,2}=\pm 3$ where the 2nd nanowire is the switching wire. We choose to satisfy $\sum_{i=1}^{n}v_{i,2}=3$. One such state is $[2,-1]$. The corresponding VSR is shown in Fig.3a. It is important to note that the states $[3,0]$, $[1,-2]$, $[0,-3]$, etc. also satisfy $\sum_{i=1}^{n}v_{i,2}=3$ while also making the middle nanowire switch first as the bias current of the devices is increased from zero in the positive direction. We show this for the state $v_{i}=[1,-2]$ with a simple calculation: $\sum_{i=1}^{3}v_{i,m}=v_{1,2}+v_{2,2}+v_{3,2}=1+0+(-1)(-2)=3$. Recall that $v_{i,m}=-v_{m,i}$ by definition.
@@
A b-invariant PSD can also be realized from systems with identical superconducting nanowires that are not equidistant. To illustrate this, we did the following: (1) take an array of two rational numbers $[2/3,5/6]$; (2) appended 0 and 1 to this array $[0,2/3,5/6,1]$; (3) computed the average of this array ($5/8$); then (4) placed a 5th nanowire at this array at location $5/8$, so the final locations of the nanowires are $[0,5/8,2/3,5/6,1]$. This array exemplifies an asymmetrical SQUID. For all wires, choose $\phi_{c}=12\pi/5$, so that $n\phi_{c}/2\pi=6$ is an integer, where $n=5$. Then, based on Eq.(7), get $\sum_{i=1}^{n}v_{i,2}=\pm 6$, choose the positive sign. Then, the vorticity state $[3,-1,0,0]$ that satisfies this constraint. Indeed, $\sum{v_{i,m}}=v_{1,2}+v_{2,2}+v_{2,3}+v_{2,4}+v_{2,5}=3+0+-(-1)+-(-1)+-(-1)=6$. The 2nd wire is the switching wire (assuming the bias current is ramped up). The VSR corresponding to this state is shown in Fig. 3b. This is a b-invariant PSD.
@@
The PSDs examples considered above require an exact tuning of the nanowire properties or the magnetic field. However, it turns out that a small disorder in the nanowire location and/or their critical phase converts PSD into a superconducting range controller (SRC). This can be illustrated by introducing disorder into the parameters of the nanowires in a symmetrical 3-SQUID, transforming Fig.3a to Fig.2b. The VSR triangle is tilted due to the disorder. Recall that the boundaries of the vorticity stability region (VSR) represent the critical currents of the device. Here we observe that the right side of each triangle is lifted. The right triangle (blue) is shifted completely into the positive current domain. Thus the current must be larger than some above-zero threshold in order to flow without any voltage applied. This is the essence of the SRC. Such device permits a supercurrent in a specific range, $[I_{c,-}>0,I_{c,+}>0]$. This means that the device is not superconducting if there is no supercurrent applied to it. This SRC effect is stable against small variations of the device parameters.

### v3

@@ Abstract
Diodes have a nonreciprocal voltage versus current relationship, produced by breaking the space and time reversal symmetry. However, developing high-end superconducting computers requires a superconducting analogue of the traditional semiconductor diode. Such a superconducting diode exhibits non-reciprocity, or a high asymmetry in its critical currents. We present a model of a perfect superconducting diode based on a superconducting quantum interference device made with multiple superconducting nanowires. The diode predicted by our model has a large positive critical current, while the negative critical current can be exactly zero. This 100% diode efficiency ($\eta=1$) remains stable against small changes of the magnetic field. Another important result is that under certain and quite broad conditions such devices can act as supercurrent range controllers. In such device a supercurrent can flow with zero voltage applied, but only if the supercurrent is contained in some narrow, adjustable range, which excludes zero current.
@@ I Introduction
Superconducting electronics offer a significant energy advantage over that of their traditional semiconductor counterpart due the absence of resistivity Tinkham (2004); Bezryadin (2012). This drastically improves energy scalability of superconducting computing resources.
@@
Superconducting diodes are useful for quantum and cryogenic electronics with their ability to allow supercurrent to flow preferentially in one direction Souto, Leijnse, and Schrade (2022); Chen et al. (2018); Hoshino et al. (2018); Baumgartner et al. (2021). In quantum computing, superconducting diodes are crucial for efficient power delivery in quantum processors, ensuring low-noise operation at ultra-cold temperatures Ingla-Aynés et al. (2025); Ideue and Iwasa (2020). In cryogenic electronics, superconducting diodes enable rectification circuits that convert AC to DC in cryogenic temperatures which improves energy efficiency Tuhin (2025). Unique properties of superconducting diodes make them ideal for ultra-sensitive detection systems and sensors used in medical imaging Ando et al. (2020); Hou et al. (2023). Superconducting diodes might play a role in advanced superconducting circuits and enable the development of the next-generation superconducting logic devices and memory elements Ma, Zhan, and Lin (2025); Alam et al. (2021); Ilin et al. (2021); Song et al. (2023). Note that an intrinsic perfect superconducting diode (PSD) has been discussed theoretically Chakraborty and Black-Schaffer (2024). Additionally, Josephson Junctions have been used, but are limited due to their large size Holmes (2021); Herr, Josephsen, and Herr (2023). Therefore, new approaches are necessaryGolod, Morlet-Decarnin, and Krasnov (2023). We propose an alternative approach to construct PSD using one-dimensional superconducting nanowires which doesn’t require any exotic superconductor and can be made out of ordinary superconducting metals such as niobium or aluminum.
@@
We present a simple yet efficient model of a superconducting quantum interference device (SQUID) made with a few one-dimensional nanowires acting as parallel weak links. One key difference with respect to the ordinary SQUID is that our nanowires are assumed to have a linear current-phase relationship, as established previouslyMurphy and Bezryadin (2017); Murphy, Averin, and Bezryadin (2017); Ilin et al. (2021); Bezryadin (2012). The phase difference $\phi$ is limited by the critical phase $\phi_{c}$, proportional to the wire length. The linear CPR means that the supercurrent, $I$, is proportional to the phase gradient of the condensate wave function (similar to free quantum particle). The phase gradient, on the other hand, is proportional to the phase difference $\phi$ between the ends of the nanowire. The linear CPR enables the manifestation of multiple highly stable vorticity states. Our modelSun and Bezryadin (2025, 2024) allows us to calculate vorticity stability regions (VSR), which are the regions in the current-magnetic (I-B) field plane within which the total supercurrent is within the negative and positive critical current. We discover that some of the vorticity states generate VSRs that act as an supercurrent range controller (SRC) or as a perfect superconducting diode (PSD). In an SRC, the supercurrent ($I$) is confined between two positive values, i.e. $0<I_{c,-}<I<I_{c,+}$. In a PSD, the supercurrent is confined as $0=I_{c,-}<I<I_{c,+}$, which corresponds to the perfect efficiency $\eta=(I_{c,+}+I_{c,-})/(I_{c,+}-I_{c,-})=1$. Note that since in the SRC $I_{c,-}>0$, therefore formally speaking $\eta>1$.
@@
II Linear CPR model for a multiple wire SQUID
We consider a parallel array of superconducting nanowires connected to two macroscopic thin-film superconducting electrodes. To model this device, we assume that all the superconducting nanowires obey a linearTinkham (2004) current-phase-relationship (CPR) as given by:
@@ II Linear CPR model for a multiple wire SQUID
where $i$ is the wire number, $\phi_{i}$ is the phase difference across the wire $i$, and $\phi_{c,i}$ and and $I_{c,i}$ represent the critical phase and the critical current of the wire $i$. In other words, if the current in the wire $i$ equals its critical current, $I_{c,i}$, then the phase across the wire is defined as the critical phase, $\phi_{c,i}$, of the wire. The phase across the wire (or the phase bias of the wire) is defined as the phase difference of the superconducting condensate complex wave function, taken between the ends of the wire. The ends of the wire are the points where the wire connects to macroscopic superconducting electrodes.
III Switching current criterion
In this section we elucidate at which point our multi-wire superconducting quantum interference device (MW-SQUID) switches to the normal state. To confirm our conclusions we add experimental results, which illustrates that our choice for the critical current criterion for the entire device is reasonable.
The starting point for our model is the notion that each wire has a critical current $I_{c,i}$, such that if the bias current in the wire, $I_{i}$, is larger than $I_{c,i}$ then the wire switches to the resistive, normal state, i.e., it becomes non-superconducting. The key criterion of the model is that the critical current of the entire device (MW-SQUID) is reached if the current in any of the wires reaches its critical value for that wire. Suppose the bias current of the entire device is $I$, and the critical current of the device is $I_{c}(v_{k})$, defined for a given fixed distribution of vortices $v_{k}$ in the cells between the wires. Here $k$ (integer) is the cell number. With these notations, the critical current criterion of our model is $I=I_{c}(v_{k})$ if $I_{i}=I_{c,i}$ for any wire number $i$. Since the vorticity, $v_{k}$, can accept different values, one expects that the critical current will be a multi-valued function of the magnetic field. This is confirmed experimentally as is shown below.
Here we discuss in detail the critical current criterion, stating that $I=I_{c}(v_{k})$ if $I_{i}=I_{c,i}$. The key idea is that the critical current function, $I_{c}(v_{k})$, for a given vorticity state represents the boundary of the corresponding vorticity stability region (VSR). To understand this, consider the case when the current is ramped up from zero. At some time the current in one wires becomes equal to the critical current of this wire $i$, $I_{i}=I_{c,i}$. Then the wire cannot remain superconducting. Then one of two scenarios might happen. (1) If there is some other vorticity state in which the currents in all wires are subcritical (for the given total bias current $I$) then there will be a phase slip in the wire which case the numbers of vortices trapped between the wires to change in such a way as to reduce the current in the wire $i$. As a result of such phase slip some heat is generated and the entire device typically also switches to a non-superconducting regime, since its temperature is increased. (2) If the bias current is low then the heat generated might be insufficient to switch the entire device. Then the event of vortex redistribution is classified as a hidden phase slip. Such hidden phase slips can be observed by using a special technique reported in Murphy et al.Murphy and Bezryadin (2017). In both cases, the current vorticity state has been changed or destroyed. This means that the boundaries of the VSR is the critical current for that specific vorticity.
Figure 1: 2-SQUID experimental data with model fits. The model parameter fits can be found in Table 1. This experimental data and our model fits validates our vorticity switching assumption, where if one nanowire reaches its critical current, then the vorticity of the device will switch due to a phase slip.
It is clear that a given fixed vorticity distribution is stable only as long as all wires remain subcritical ($I_{i}<I_{c,i}(v_{k})$). As soon as one of the wire reaches its critical point ($I_{i}=I_{c,i}(v_{k})$) the vorticity state must change. Such reshuffling of vortices might cause the device to switch to the normal state with some probability. To compare with experimental results, we calculate critical currents for many vorticity states and thus obtain a multi-valued function $I_{c}(v_{k},B)$. Then some branches of this function can be observed to correspond to the switching events of the entire device. However, some branches might not manifest themselves under typical experimental conditions.
We validated this switching assumption by fabricating a simple 2-SQUID, i.e., a SQUID with two superconducting nanobridges, parallel to each other and connecting larger electrodes. We measure the critical current of the device versus perpendicular magnetic field, and we performing fits using our model. The 2-SQUID sample was fabricated on Si substrate covered with oxide $SiO_{2}$ and a low stress $SiN$ on top. A trench with two bridges was fabricated using standard electron beam lithography and reactive ion etching. A hydrofluoric acid etch selectively removed the underlying $SiO_{2}$, creating an undercut that fully suspends the $SiN$ nanobridges. Photolithography was subsequently employed to define electrode patterns, followed by $Nb$ deposition via sputtering and a lift-off process to complete the device structures. A zoomed out SEM image of the sample can be found in Fig. 2a. A zoomed in image on one of the bridges is shown in Fig. 2b. Our measurement set-up is illustrated in Fig. 2c. It contains a controllable voltage source and a resistor connected in series with the sample to realized a current biasing experiment. Voltage is measured on the series resistor (to calculate the current by Ohm’s law) and on the sample. The voltage-current curves are then plotted in LabView. They normally show a voltage jump at some current, which is then taken as the critical current of the entire device.
It is important to note that we have normalized the critical currents such that $j_{c,i}=I_{c,i}/\langle I_{c,i}\rangle$ where $\langle I_{c,i}\rangle$ is the average of all the critical currents, $\langle I_{c,i}\rangle=\sum_{i}I_{c,i}/n$. This is so $\sum_{i}j_{c,i}=n\langle I_{c,i}\rangle/\langle I_{c,i}\rangle=n$ where $n$ is the number of wires in the SQUID. In the example considered in this section $n=2$. Therefore, the maximum normalized critical current of the entire device equals $2$ .
$\phi_{c,1}$ (radians) $\phi_{c,2}$ (radians) $j_{c,1}$ $j_{c,2}$ $\Delta B$ (Gauss)
70 70 1.05 0.97 0.027
Table 1: Model Fitting Parameters
The measured critical current curves are shown in Fig. 1. The positive critical current is shown by blue triangles and the negative critical current is shown by red dots. The positive and negative critical currents were normalized by dividing by half of the maximum critical current($I_{max}=1.18mA)$, corresponding to the average value of all measured critical currents. The experimental magnetic field was then referenced to zero field by exploiting the intrinsic symmetry of the $I_{c}$$B$ pattern and shifting the $x$-axis such that $B=0$ satisfied this symmetry. Finally, the field axis was scaled by $\delta B$ to align the experimental data with the model.The function of the critical current plotted versus magnetic field appears multi-valued. This fact reflects the expectation that different vorticity states have different critical currents. The corresponding fits produced by our model are shown by blue and red lines (Fig. 1). Note, the experimental data and the model’s total critical current were normalized. The data fit demonstrates strong evidence that our model assumption, the criterion for the switching of the device, is a good description of such nanowire SQUIDs.
In Fig. 1, the experimental data branches corresponds to the right sides of the diamond VSRs. Each branch corresponds to a unique vorticity state. Our switching assumption states that if the critical current of the VSR is reached, then the device will switch its vorticity if and only if there is a vorticity state for which the current in the critical wire is less than its critical current (for a given magnetic field). On the contrary, if $I_{c}(v_{k},B)>I_{c}(v^{{}^{\prime}}_{k},B)$, for any other vorticity distribution $v^{{}^{\prime}}_{k}$ then the device won’t switch its vorticity state, but rather turn normal. But if $I_{c}(v_{k},B)<I_{c}(v^{{}^{\prime}}_{k+1},B)$, then the device switches to $v^{{}^{\prime}}_{k}$. With a nonzero probability such vorticity change cause enough dissipation and switch the device to the normal state, especially if the current is near the critical current. The fact that we observe a multi-valued function (after many repetitive measurements of the critical current gives us evidence that our interpretation that a single phase slip can switch the device is correct. Moreover, the observed agreement between the fits and the data confirms our model.
[…]
IV Effect of the magnetic field and vortices on the critical current
@@
Using Eqs.(3),(4) we generate vorticity stability regions (VSR), each defined as region on the IBI-B plane where for a fixed vorticity, the current in each wire is lower than its corresponding critical current. The boundary of the VSR is the critical current of the device as a function of the magnetic field. If either the total current, $I$, or the magnetic field, $B$ is changed beyond the VSR boundaries, then the device switches to the normal (i.e., resistive) state.
V Results
@@
Figure 4: A SRC is demonstrated for VSRs generated by (a) a 2-SQUID with critical phases set to $\pi,2\pi]$ and critical currents set to $[2/3,4/3]$. The plotted vortices are 1-1 (red), 0 (blue), and 11 (black). And (b) a disordered 3-SQUID with wires at $[0,0.45,1]$, critical phases set to $[6,7,5]$, and critical currents set to $[1.01,1.12,0.87]$. The vortices from left to right are [2,1][2,-1] (red), $[3,0]$ (black), $[4,1]$ (blue). The grey lines indicate any other VSRs corresponding to different vorticity states.
Let us first construct an example of a point-wise PSD. Suppose the SQUID has only two wires, $n=2$, with $v_{1,2}$ vortices trapped between them (see Fig.3). Then the MPC equation is $\phi_{2}=\phi_{1}+2\pi b-2\pi v_{1,2}$. Consider the case $v_{1,2}=0$, so $\phi_{2}=\phi_{1}+2\pi b$. Zero critical current means that even if the bias current is zero the magnitude of the current in one of the wire is equal to its critical current. Therefore, either $\phi_{1}=\phi_{c,1}$ or $\phi_{1}=-\phi_{c,1}$. Suppose $\phi_{1}=-\phi_{c,1}$. Then $\phi_{2}=-\phi_{c,1}+2\pi b$. Under such condition one can increase the bias current and the device will remain superconducting. But a decrease of the bias current is not possible since the first wire would be pushed into the normal state. To construct a PSD, require that the total current is zero, $I=0$, and that the first wire is in the critical state. If $I=0$ then the current in the first wire is equal in magnitude and opposite in sign to the current in the second wire, i.e. $\phi_{1}I_{c,1}/\phi_{c,1}=-\phi_{2}I_{c,2}/\phi_{c,2}$. If $\phi_{1}=-\phi_{c,1}$ then $\phi_{c,1}I_{c,1}/\phi_{c,1}=-(-\phi_{c,1}+2\pi b)I_{c,2}/\phi_{c,2}$. So $I_{c,1}=(-\phi_{c,1}+2\pi b)I_{c,2}/\phi_{c,2}$. For this to be true the right hand side needs to be positive, because $I_{c,1}$ represents the magnitude of the critical current of the first wire. As an example, choose $b=1$ and $\phi_{c,1}=\pi$. Then $I_{c,1}=\pi I_{c,2}/\phi_{c,2}$. Suppose $\phi_{c,2}=2\pi$. Then $I_{c,1}=I_{c,2}/2$. This argument indicates that a 2-SQUID with critical phases $\pi,2\pi]$ and critical currents, for example, $[2/3,4/3]$ should have a zero critical current at the point $b=1$, i.e., it should act as a perfect superconducting diode. This device will be called point-wise PSD since $I_{c,-}(b)=0$ is only true if $b=1$, or other integers if $v_{1,2}\neq 0$ (Fig.4a).
If $v_{1,2}=0$ and if $b>1$, then the allowed range of the supercurrent becomes strictly above zero (Fig.4a). For example, if $b=1.25$ then the allowed range of the supercurrent is $0.3<I<1$. Thus at non-zero magnetic fields our PSD becomes a supercurrent range controller (SRC). The VSR repeats with the period $\Delta b=1$, due to the Little-Parks effect in this 2-SQUID.
@@
For the critical current of one polarity to be independent of the magnetic field, it must be the case that the $b$ term drops out. In other words, $2\pi b\sum_{i=1}^{n}(x_{i}-x_{m})=0$ or:
@@
Figure 5: A b-invariant PSD can be constructed by (a) a symmetrical 3-SQUID with $\phi_{c}=2\pi$ for all wires, setting vorticity states [1,-2] (red) and [2, -1] (blue) and by (b) a non-symmetrical 5-SQUID with wires at $[0,5/8,2/3,5/6,1]$ and critical phases set to $12\pi/5$. The vorticity state is [3,1,0,0][3,-1,0,0]. In both figures, the underlying gray lines indicate other VSRs corresponding to different vorticity states.
@@
Next, we construct an example of a b-invariant PSD in a symmetrical 3-SQUID. Consider a 3-SQUID device with all critical phases set to $2\pi$. According to Eq.(7), to produce a VSR with the flat and horizontal bottom one needs $\sum_{i=1}^{n}v_{i,2}=\pm 3$ where the 2nd nanowire is the switching wire. We choose to satisfy $\sum_{i=1}^{n}v_{i,2}=3$. One such state is [2,1][2,-1]. The corresponding VSR is shown in Fig.5a. It is important to note that the states $[3,0]$, [1,2][1,-2], [0,3][0,-3], etc. also satisfy $\sum_{i=1}^{n}v_{i,2}=3$ while also making the middle nanowire switch first as the bias current of the devices is increased from zero in the positive direction. We show this for the state $v_{i}=[1,-2]$ with a simple calculation: $\sum_{i=1}^{3}v_{i,m}=v_{1,2}+v_{2,2}+v_{3,2}=1+0+(-1)(-2)=3$. Recall that $v_{i,m}=-v_{m,i}$ by definition.
@@
A b-invariant PSD can also be realized from systems with identical superconducting nanowires that are not equidistant. To illustrate this, we did the following: (1) take an array of two rational numbers $[2/3,5/6]$; (2) appended 0 and 1 to this array $[0,2/3,5/6,1]$; (3) computed the average of this array ($5/8$); then (4) placed a 5th nanowire at this array at location $5/8$, so the final locations of the nanowires are $[0,5/8,2/3,5/6,1]$. This array exemplifies an asymmetrical SQUID. For all wires, choose $\phi_{c}=12\pi/5$, so that $n\phi_{c}/2\pi=6$ is an integer, where $n=5$. Then, based on Eq.(7), get $\sum_{i=1}^{n}v_{i,2}=\pm 6$, choose the positive sign. Then, the vorticity state [3,1,0,0][3,-1,0,0] that satisfies this constraint. Indeed, $\sum{v_{i,m}}=v_{1,2}+v_{2,2}+v_{2,3}+v_{2,4}+v_{2,5}=3+0+-(-1)+-(-1)+-(-1)=6$. The 2nd wire is the switching wire (assuming the bias current is ramped up). The VSR corresponding to this state is shown in Fig. 5b. This is a b-invariant PSD.
@@
The PSDs examples considered above require an exact tuning of the nanowire properties or the magnetic field. However, it turns out that a small disorder in the nanowire location and/or their critical phase converts PSD into a superconducting range controller (SRC). This can be illustrated by introducing disorder into the parameters of the nanowires in a symmetrical 3-SQUID, transforming Fig.5a to Fig.4b. The VSR triangle is tilted due to the disorder. Recall that the boundaries of the vorticity stability region (VSR) represent the critical currents of the device. Here we observe that the right side of each triangle is lifted. The right triangle (blue) is shifted completely into the positive current domain. Thus the current must be larger than some above-zero threshold in order to flow without any voltage applied. This is the essence of the SRC. Such device permits a supercurrent in a specific range, $I_{c,-}>0,I_{c,+}>0]$. This means that the device is not superconducting if there is no supercurrent applied to it. This SRC effect is stable against small variations of the device parameters.
VI Discussion